Microsoft has 890 million shares that currently trade at $399. You own 100 shares.

Part (1)

What is the total value of your shares?

Part (2)

If Apple does a 3-for-1 stock split, what will be the new share price?

Part (3)

What is the total value of your shares after the stock split?

Part (4)

If Apple pays a 20% stock dividend instead of the stock split, what will be the new share price?

Part (5)

What is the total value of your shares after the stock dividend?
Answer #1

1.

Total Value = 399(100) = $39,900

2.

New Share Price = 399/3 = $133

3.

Total Value = 133(300) = $39,900

4.

New Share Price = 399/1.2 = $332.50